A sample is analyzed to determine the concentration of an analyte. Under the conditions of the analysis the sensitivity is 17.2 ppm-1. What is the analyte’s concentration if S total is 35.2 and S reag is 0.6?


1
Expert's answer
2022-05-24T02:58:02-0400

Sensitivity (k) = 17.2 ppm-1 Stotal = 35.2 Sreag =0.6 Stotal = Average Signal (kCA) + Sreag , where CA is concentration therefore CA=(Stotal -Sreag)/k=(35.2-0.6)/17.2=2.01 ppm